write the equation of the line that passes through the point (1,1) and is parallel to the graph of y=2x+6. then graph both lines. i dont get this, help???
Answer by mananth(12270)

y = 2 x + 6
Compare this equation with y=mx+b
slope m = 2
The slope of a line parallel to the above line will be the same
The slope of the required line will be 2
m= 2 ,point ( 1 , 1 )
Find b by plugging the values of m & the point in y=mx+b
1 = 2.00 + b
b= -1
m= 2
Plug value of the slope and b in y = mx +b
The required equation is y = 2 x -1
y = 2 x -1
m.ananth@hotmail.ca